OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y -3x - 2?

OpenStudy (anonymous):

this is after the y

OpenStudy (austinl):

Ok, what are the ordered pairs that you were referring to?

OpenStudy (anonymous):

(0, -1) (0, 0) (-1, 0) (1, 1)

OpenStudy (austinl):

Ok, this is fairly simple. Ordered pairs look like this (x,y) What you would do is take the ordered pairs and plug the x and y values into the inequality. If the pair make the inequality true, then it is a solution. Make sense?

OpenStudy (anonymous):

kinda, would it be the 3rd answer choice ?

OpenStudy (austinl):

\[y \ge -3x-2\] \[-1 \ge -3(0) - 2\] \[-1 \ge -2\] That is true, so then yes!

OpenStudy (anonymous):

yay ! thank you so much !

OpenStudy (austinl):

no problem! it was my pleasure.